A die is thrown 150 times and getting 1 or 3 is considered success. Find the mean, variance of number of successes

Grade:Upto college level

1 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student,
Please find the answer to your question below
From given data
A die is thrown 150 times and getting 1 or 3 is considered success
So p=1/3 and q=2/3 and n=150
mean=np=150(1/6+1/6)=150/3=50
Variance=npq=150(1/3)(2/3)=33.34